# MapGuide: A Simple yet Effective Method to Reconstruct Continuous Language from Brain Activities

### v1

@@ 3.2 Stage A: Mapping from Brain Activities to Text Embeddings
We have developed a mapper to predict fMRI to text features while enhancing fMRI representation robustness. Our mapper consists of the fMRI encoder $\mathcal{E}$ and the text embedding projector $\mathcal{D}$. The encoder $\mathcal{E}$ processes fMRI data into a latent space, which the text embedding projector then translates into text embeddings. We first optimize the Mean Squared Error (MSE) loss between predicted and ground truth text embeddings. The loss $\mathcal{L}_{MSE}$ is formulated as:
@@
To learn denoised fMRI representations, our model further incorporates contrastive learning with random masking technique $M(\cdot,\text{ratio})$, generating masked data $F^{m}=M(F,\text{ratio})$ and treating masked samples as positive samples to the unmasked input. The fMRI projector $\mathcal{P}$ derives the hidden layer representation, with the InfoNCE loss calculated as follows:
@@ 4.2 Dataset
We use the dataset provided by LeBel et al. (2023) to evaluate decoding performance, concentrating on perceptual speech task responses. The dataset includes a set of training stories and one test story, comprising 27,449 fMRI samples in the training set and 291 in the test set, collected from three subjects who listened to identical training and testing stories. The fMRI data were acquired with a TR of 2 seconds. The dataset also includes word information and timestamps for each word. We selected 10,000 cortical voxels for decoding purposes, consistent with those chosen by Tang et al..
@@ 4.4 Baseline
In our study, we have chosen the approach by Tang et al. as the baseline for text decoding. This decision was influenced by two key factors. Firstly, our decoding and evaluation process is uniquely designed to be stepwise and timestamp-based. Secondly, the multi-subject model utilized in studies such as Xi et al. (2023) significantly differs from the single-subject focus of our task, making them less suitable for direct comparison.
@@ 4.5 Implementation
To implement Stage A, we employ the Huggingface Transformers library based on PyTorch. While training the models in Stage A, we partition the datasets into training and validation sets with an 80% - 20% split ratio. We train separate models for each of the three subjects. For Stage B, we use the same structure of our baseline. Since changing the hyperparameters of Stage B will affect the random generation, we use the same hyperparameters by Tang’s method. Additionally, we leverage the pre-trained GPT model introduced by Tang for two purposes: feature extraction and text generation. All experimental procedures are carried out using NVIDIA GeForce GTX 1080 Ti GPUs. The hyper-parameter settings to achieve the best performance are detailed in the ablation study in Section 5.2.
Figure 2: An error band line chart depicting the relationship between the embedding prediction performance of Stage A’s learned mapper and the final text reconstruction metrics. We plot with the BERT and METEOR scores on all three subjects.
@@ 5 Results
In the results section, we will first compare the metrics of text reconstruction of MapGuide against the previous SOTA. We will then conduct a detailed ablation study to evaluate the efficacy of MapGuide’s modules and discuss the influence of hyper-parameter setting. We will lastly do a correlation analysis of the experimental results.
@@ 5.1 Reconstruction Results
To have a qualitative impression of the text reconstructed by our model, we randomly selected some samples and depicted them in Table 2. As shown in the samples, though not fully resembling the ground truth, the texts reconstructed by our model have several fragments semantically identical or similar to the ground truth.
@@ 5.3 Correlation Analysis
In previous sections, we observed a tendency that the better the mapper performs in Stage A to predict text embeddings, the more likely the mapper can guide a better text reconstruction in Stage B. This is also intuitive since a high-quality mapper could more accurately guide the text generator to reconstruct semantic-related contents. In this section, we will check whether such intuitions comply with our experimental results.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments. This research was supported by grants from the National Natural Science Foundation of China to S. W. (62036001) and S.W. (the STI2030-Major Project, grant number: 2021ZD0204105).

### v2

@@ 3.2 Stage A: Mapping from Brain Activities to Text Embeddings
We have developed a mapper to predict fMRI to text features (see Appendix B for details) while enhancing fMRI representation robustness. Our mapper consists of the fMRI encoder $\mathcal{E}$ and the text embedding projector $\mathcal{D}$. The encoder $\mathcal{E}$ processes fMRI data into a latent space, which the text embedding projector then translates into text embeddings. We first optimize the Mean Squared Error (MSE) loss between predicted and ground truth text embeddings. The loss $\mathcal{L}_{MSE}$ is formulated as:
@@
To learn denoised fMRI representations, our model further incorporates contrastive learning with random masking technique $M(\cdot,\text{ratio})$, generating masked data $F^{m}=M(F,\text{ratio})$ and treating masked samples as positive samples to the unmasked input. Without loss of generality, we use infoNCE loss as the loss function for contrastive learningOord et al. (2019). The fMRI projector $\mathcal{P}$ derives the hidden layer representation, with the infoNCE loss calculated as follows:
@@ 4.2 Dataset
We use the dataset provided by LeBel et al. (2023) to evaluate decoding performance, concentrating on perceptual speech task responses. The dataset includes a set of training stories and one test story, comprising 27,449 fMRI samples in the training set and 291 in the test set, collected from three subjects who listened to identical training and testing stories. The fMRI data were acquired with a TR of 2 seconds. The dataset also includes word information and timestamps for each word. We select 10,000 cortical voxels for decoding purposes, consistent with those chosen by Tang et al..
@@ 4.4 Baseline
In our study, we have chosen the approach by Tang et al. as the baseline for text decoding. This decision was influenced by two key factors. Firstly, our decoding and evaluation process is uniquely designed to be stepwise and timestamp-based. This choice is based on the practical application scenarios of neural decoding, particularly in the field of brain-computer interfaces. For practical human needs, the ability to collect signals step-by-step during human language expression and generate words accordingly is more in line with the application scenario of human conversation. Therefore, starting from practical application scenarios, we prefer to choose a task form that is closer to real-world use. Secondly, the multi-subject model utilized in studies such as Xi et al. (2023) significantly differs from the single-subject focus of our task, making them less suitable for direct comparison.
@@ 4.5 Implementation
To implement Stage A, we employ the Huggingface Transformers library based on PyTorch. While training the models in Stage A, we partition the datasets into training and validation sets with an 80% - 20% split ratio. We train separate models for each of the three subjects. For Stage B, we use the same structure of our baseline. Since changing the hyperparameters of Stage B will affect the random generation, we use the same hyperparameters by Tang’s method. Additionally, we leverage the pre-trained GPT model introduced by Tang for two purposes: feature extraction and text generation. All experimental procedures are carried out using eight NVIDIA GeForce GTX 1080 Ti GPUs. The hyper-parameter settings to achieve the best performance are detailed in the ablation study in Section 5.2.
Figure 2: An error band line chart depicting the relationship between the embedding prediction performance of Stage A’s learned mapper and the final text reconstruction metrics. We plot with the BERT and METEOR scores on all three subjects, with the statistical scope covering all models mentioned in the ablation experiments and the optimal upper-bound. The values of embedding prediction performance have been standardized.
@@ 5 Results
In this section, we will first compare the metrics of text reconstruction of MapGuide against the previous SOTA. We will then conduct a detailed ablation study to evaluate the efficacy of MapGuide’s modules and discuss the influence of hyper-parameter setting. We will lastly do a correlation analysis of the experimental results.
@@ 5.1 Reconstruction Results
To have a qualitative impression of the text reconstructed by our model, we randomly select some samples and depict them in Table 2. As shown in the samples, though not fully resembling the ground truth, the texts reconstructed by our model have several fragments semantically identical or similar to the ground truth.
@@ 5.3 Correlation Analysis
In previous sections, we observe a tendency that the better the mapper performs in Stage A to predict text embeddings, the more likely the mapper can guide a better text reconstruction in Stage B. This is also intuitive since a high-quality mapper could more accurately guide the text generator to reconstruct semantic-related contents. In this section, we will check whether such intuitions comply with our experimental results.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments.
This research was supported by grants from the National Natural Science Foundation of China to S. W. (62036001) and S.W. (the STI2030-Major Project, grant number: 2021ZD0204105).
@@ Appendix A Details of Metrics
Appendix B Acquisition of Text Embeddings
We replicate the methodology outlined in prior research by Tang et al. to generate a stimulus matrix that corresponds to the fMRI data. For each word-time pair $(s_{i},t_{i})$ within every narrative, we input the word sequence $(s_{i-5},s_{i-4},\ldots,s_{i-1},s_{i})$ into a language model. From the model’s hidden layer, we extract semantic features of $s_{i}$, resulting in a revised list of vector-time pairs $(M_{i},t_{i})$, where $M_{i}$ signifies an $n$-dimensional semantic embedding for $s_{i}$. These pairs are resampled utilizing a three-lobe Lanczos filter to synchronize the vectors with the fMRI acquisitions. Subsequently, we employ a linearized Finite Impulse Response (FIR) model to fit every cortical voxel in each subject’s brainHuth et al. (2016). For each of the $n$ features, we apply a distinct linear temporal filter with four delays ($t-1$, $t-2$, $t-3$, and $t-4$ timepoints), resulting in a total of $4n$ features. All punctuation is removed during the representation acquisition and text generation process.